OpenStudy (k.binks):

Solve 4x + 3y = 13 by the substitution method. x + y = 4 I understand how to substitute, but I'm unsure of the negatives in this one

rishavraj (rishavraj):

x + y = 4 x = 4 - y so now plug the value of "x" in 1st equation

OpenStudy (k.binks):

For some reason I did -y=x-4. Careless mistake, thank you again

rishavraj (rishavraj):

u should hav written tht y = 4 - x